OpenStudy (anonymous):

Philip uses his cellular phone only for local service. One month he paid $32.50 for 250 minutes of service, and another month he paid $37.00 for 400 minutes of service. Assume a linear relationship exists between the cost and the number of minutes of service. How much will Philip pay for 500 minutes of service?

OpenStudy (bahrom7893):

these are just two points: x1 = 250, y1 = 32.50 x2 = 400, y2 = 37.00 Find the equation of the line

OpenStudy (bahrom7893):

M = 5.50/150

OpenStudy (anonymous):

Well done bahrom!

OpenStudy (bahrom7893):

or M = 55/1500 = 11/300

OpenStudy (bahrom7893):

now just use any of the points: y - 37 = (11/300)(x-400). Plug in x = 500 y-37 = (11/300)*(500-400) y-37 = (11/300)*100 y-37 = 11/3 y = 11/3 + 37

OpenStudy (bahrom7893):

y = 40.67